OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later.

OpenStudy (anonymous):

6 hours later from when? y = 65 - 2x = 65 - 2*6 = 53